# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v4

@@ C.2.2 Geometric Regularization Losses: Alignment and Uniformity
where $\mathcal{F}(\rho_{1},\rho_{2})$ is the fidelity between states $\rho_{1}$ and $\rho_{2}$. Lower values of $1-\mathcal{F}$ indicate higher similarity between quantum states. If $\rho_{i}=|\psi_{i}\rangle\langle\psi_{i}|$ are pure states derived from normalized vectors $\mathbf{z}_{i}$, then $\mathcal{F}(\rho_{1},\rho_{2})=|\langle\psi_{1}|\psi_{2}\rangle|^{2}=\text{cos}^{2}(\theta)$. Therefore, alignment loss becomes:
@@
C.2.3 Overall QRGCL Objective
@@
Appendix D Details of Benchmark Models
@@
D.1 Classical RGCL (CRGCL)
@@ D.4 Quantum GNN (QGNN)
D.5 Equivariant Quantum GNN (EQGNN)
@@ D.6 QCL
D.7 Classical-Quantum CL (CQCL)
@@
D.8 Quantum-Classical GCL (QCGCL)
@@
Appendix E Details of Experimental Setup
E.1 Simulation Tools and Environment
@@ E.2 Hyperparameters and Configurations
Appendix F Details of Results and Discussion

### v5

@@ D.2.2 Geometric Regularization Losses: Alignment and Uniformity
where $\mathcal{F}(\rho_{1},\rho_{2})$ is the fidelity between states $\rho_{1}$ and $\rho_{2}$. Lower values of 11-$\mathcal{F}$ indicate higher similarity between quantum states. If $\rho_{i}=|\psi_{i}\rangle\langle\psi_{i}|$ are pure states derived from normalized vectors $\mathbf{z}_{i}$, then $\mathcal{F}(\rho_{1},\rho_{2})=|\langle\psi_{1}|\psi_{2}\rangle|^{2}=\text{cos}^{2}(\theta)$. Therefore, alignment loss becomes:
@@
D.2.3 Overall QRGCL Objective
@@
Appendix E Details of Benchmark Models
@@
E.1 Classical RGCL (CRGCL)
@@ E.4 Quantum GNN (QGNN)
E.5 Equivariant Quantum GNN (EQGNN)
@@ E.6 QCL
E.7 Classical-Quantum CL (CQCL)
@@
E.8 Quantum-Classical GCL (QCGCL)
@@
Appendix F Details of Experimental Setup
F.1 Simulation Tools and Environment
@@ F.2 Hyperparameters and Configurations
Appendix G Details of Results and Discussion